OpenStudy (anonymous):

If f(x)=2−x^1/2 and g(x)=x^2−9, what is the domain of g(x)÷f(x)?

OpenStudy (anonymous):

(−∞,2] and [2,∞) (−∞,4) and (4,∞) (−∞,2] and [4,∞) (−∞,−2] and [2,∞)

OpenStudy (solomonzelman):

Can you first write the `g(x)÷f(x)` for me please?

OpenStudy (anonymous):

x^2-9÷2-x^1/2

OpenStudy (solomonzelman):

\(\large\color{red}{ \displaystyle f(x)=2−x^{1/2} }\) \(\large\color{blue}{ \displaystyle g(x)=x^2- 9 }\) \(\large\color{black}{ \displaystyle g(x)\div f(x)=\frac{\color{blue}{ \displaystyle x^2- 9 } }{\color{red}{ \displaystyle 2−x^{1/2} }} }\)

OpenStudy (solomonzelman):

YOu probably know that: \(\large\color{black}{ \displaystyle x^{1/2}=\sqrt{x} }\)

OpenStudy (solomonzelman):

Can you take the square root of a number that is smaller than zero?

OpenStudy (solomonzelman):

I am guessing that my interpretation is incorrect... can you write the equation in code or draw to make it clear?

OpenStudy (anonymous):

You are correct i got it from here thank you so much for your help !

OpenStudy (solomonzelman):

Whatever you say... yw then